Determine the yearly periodic deposit needed to reach a financial goal of $150,000 at a rate of 5% annually for a time of 18 years?
thank you!

Answer by rfer(10432)

P=(FV*i)/(1+i)^n-1
P=(150000*.05)/((1+.05)^18)-1
P=(7500/2.41)-1
P=3112.03-1
p=$3111.03